# MemAscend: System Memory Optimization for SSD-Offloaded LLM Fine-Tuning

### v1

@@ Abstract
Owing to the huge success of generative artificial intelligence (AI), large language models (LLMs) have emerged as a core subclass, underpinning applications such as question answering, text generation, and code completion. While fine-tuning these models on domain-specific data can yield significant performance gains, it also poses daunting computational challenges, especially for researchers and small organizations with limited hardware resources. Although SSD offloading (i.e., ZeRO-Infinity) has emerged as a viable strategy to overcome the GPU memory barrier via leveraging both system memory (i.e., CPU DRAM) and storage space (i.e., solid-state devices, SSDs), its design primarily targets model-centric performance issues. As a result, key system-level issues, including system memory fragmentation, inefficient pinned buffer allocation, peak CPU usage spikes, and file system overhead, remain unaddressed, stifling scalability and inflating costs. Such an observation motivates this paper to introduce MemAscend, a framework that systematically tackles the underexplored system memory bottlenecks in SSD-offloaded LLM training, with a focus on resource-constrained environments. By streamlining pinned-memory allocation, eradicating fragmentation, and mitigating peak overhead, MemAscend reclaims a substantial system memory budget, enabling larger models, longer context windows, and higher batch sizes without exceeding modest hardware limits. Across diverse LLM benchmarks, MemAscend reduces peak system-memory consumption by an average of 55.7% compared with standard SSD offloading techniques, lowering the hardware barrier for fine-tuning and unlocking new possibilities for cost-effective large-scale training on limited-resource machines.
@@ I Introduction
The growth of large-language-model (LLM) applications has significantly surged in recent years, propelled by steady advances in model capabilities [1]. LLMs have now powered conversational assistants, automatic summarization, multilingual translation, and countless other tasks. To adapt LLMs to a specific domain, fine-tuning on domain-specific datasets is a widely used approach [2] to customize pre-trained foundational models and address the nuances of a specific domain better. Although fine-tuning markedly boosts reliability and task performance, it also demands considerable computation and memory usage. Commercial leaders routinely train foundation models on massive GPU clusters. For example, Meta’s Llama 3 relied on 16,000 NVIDIA H100 GPUs [3]. By contrast, academic labs, individual researchers, and small enterprises must operate within tight hardware budgets, with limited GPU count and modest system-memory capacity. Finding cost-effective ways to fine-tune LLMs under such tight resource constraints, therefore, remains an open challenge.
@@
Broadly, LLM training consumes two memory classes: (1) static memory, which includes model weights and optimizer states, and (2) residual states, such as activations and intermediates [4]. Static memory is substantial because modern transformers contain billions of parameters, and optimizers such as Adam store both momentum and variance for each weight [5]. Activations can also dominate due to their linear growth with batch size and sequence length [6]. Several techniques have been proposed to mitigate the above memory pressure. For instance, to lower activation memory, gradient-checkpointing [7] does not store every activation during the forward pass, but instead stores only a few checkpointed activations and recomputes those missed ones during backpropagation. Works such as Liger-Kernel [8] further fuse operations (e.g., RMSNorm, SwiGLU, Cross-Entropy [9, 10, 11]) into single Triton [12] kernels, eliminating transient buffers and launch overhead. Mixed-precision and related quantization techniques [13] shrink activations by casting tensors to fp16 or bf16; this in turn enables memory-efficient kernels such as Flash-Attention [14].
@@
For static memory, quantization has been widely explored. For instance, 8-bit optimizers [15] and 4-bit optimizers [16] have shown promising results in reducing static memory consumption with minimal impact on convergence speed. On the other hand, Parameter-Efficient Fine-Tuning (PEFT) [17], especially Low-Rank Adaptation (LoRA) [18], significantly reduce the number of trainable parameters by focusing on low-rank matrices while keeping the pre-trained model frozen. Although methods like LoRA offer substantial computational and memory efficiency, they often require meticulous hyperparameter tuning, such as determining the rank and selecting which weights to train, to achieve optimal performance [19, 20]. Consequently, full-parameter training remains crucial for maximizing task-specific performance, particularly in scenarios where pre-trained models have not been exposed to domain-specific or private data. In such cases, the discrepancies between the target and pre-trained domains can be substantial, necessitating full-parameter training to bridge the gap effectively.
@@
Although full-parameter fine-tuning can achieve optimal performance in fine-tuning scenarios, the large size of pre-trained models limits the feasibility of training in resource-limited environments. Typically, fine-tuning tasks involve relatively smaller datasets, so the primary bottleneck in resource-limited scenarios is that GPU memory or system memory resources are insufficient to handle the static and residual memory requirements of these models. To address these limitations, SSD offloading has emerged as a promising solution [21]. This technique transfers static memory, such as model weights and optimizer states, from GPUs to SSDs, which generally provide significantly larger storage capacity. For example, ZeRO-Infinity, a state-of-the-art SSD offloading system, moves static memory to NVMe SSDs. In this approach, GPU memory is reserved for intermediate values, while system memory acts as an intermediary for transferring data between GPUs and SSDs or for performing operations, such as optimizer steps, on the CPU. This strategy reduces the volume of data transfer across the PCIe bus. Although SSD offloading may reduce training performance due to additional PCIe transfer latency, this impact is mitigated for fine-tuning tasks, which typically involve relatively smaller dataset compared to pre-trained task. Furthermore, SSD offloading integrates effectively with other optimization techniques such as Flash-Attention, Liger-Kernel, gradient checkpointing, and even advanced strategies to offload checkpointed values into system memory. By integrating these techniques, GPU memory usage is minimized, enabling the largest trainable context length or batch size, and potentially mitigating the original performance loss caused by additional data transfers across heterogeneous devices.
However, SSD offloading systems like ZeRO-Infinity are primarily designed to scale trainable model sizes on high-performance machines rather than hardware with limited system memory. While GPU memory usage has been extensively studied, system memory usage has received comparatively less attention. Several studies have explored SSD offloading techniques. For example, LoHan [22] enables fine-tuning of very large models on a single consumer-grade GPU with limited main memory through holistic tensor movement management. Although impressive in extreme resource constraints, its optimizations are specifically tailored for this single-GPU scenario and address a problem context that is different from that of enhancing the general distributed ZeRO-Infinity system. Other approaches, such as Smart Infinity [23], focus on specialized architectures that enable the computation on SSDs and are not broadly applicable to commodity hardware. In our study, we analyze the system memory requirements for SSD offloading and identify key challenges. System memory, acting as an intermediary, includes various types of pinned buffers [24] that facilitate data transfer between GPU and CPU. Inefficient management of these buffers leads to fragmentation and wasted memory. Additionally, significant peak memory overhead arises from inefficient operations involving the CPU and system memory. Furthermore, system memory may also accommodate activation checkpointed values when the offloaded activation checkpointing technique is employed. This technique stores only a few checkpointed activations and recomputes others when needed. By offloading these checkpointed values into system memory, it is possible to save substantial GPU memory space, thereby enabling the training of tasks with the largest context lengths. However, the original SSD offloading system introduces unnecessary memory overhead, which undermines the feasibility of the offloaded activation checkpointing method.
In this paper, we focus on two key types of system memory inefficiencies: fragmentation and reduction of peak memory overhead in SSD offloading scenarios. These inefficiencies arise primarily due to the limitations of existing SSD offloading systems for resource-limited machines, particularly the constrained resources of both GPU memory and system memory. Furthermore, standard memory management practices, such as aggressive alignment in pinned memory allocation, while beneficial in other contexts, introduce significant overhead and waste in the large, static allocation patterns typical of SSD offloading. These constraints result in considerable resource wastage and hinder the potential for cost-effective SSD offloading. To address these challenges, we propose solutions that minimize system memory requirements while maximizing the trainable model size and the trainable context length within the same configuration. The contributions of this study are summarized as follows:
@@
    The issue of buffer fragmentation was addressed, with a proposed solution achieving an average reduction of 72.71% in various models.
@@
    Pinned memory allocation overhead was significantly reduced by eliminating unnecessary alignment, saving substantial system memory.
@@
    Overhead and double-up peak memory problems during overflow checks were mitigated, eliminating the 1.25x memory overhead.
  4. 4.
    Inefficiencies in the system memory to SSD I/O pathway were resolved, with a proposed solution reducing I/O latency, particularly for write operations.
  5. 5.
    Integration of a pure half-precision optimizer into the SSD offloading system reducing I/O transfer volume by 58%, improving throughput by an average of 17-27%.
In summary, while fine-tuning LLMs is essential in many cases, most machines struggle to meet the requirements for training these models. Even during the fine-tuning phase, where the dataset is relatively small compared to the trillions of tokens used in the pre-training phase, a significant number of GPUs to maximize training throughput is often unnecessary. The primary objective and bottleneck lies in accommodating the model size and extended context lengths, which demand substantial space to store static and residual state memory. An SSD offloading system shows promise in addressing these challenges by transferring static memory to SSDs. Additionally, integrating this system with techniques such as Liger-Kernel, offloaded activation checkpointing, and gradient accumulation can reduce activation memory usage and improve throughput. However, current SSD offloading systems do not efficiently manage system memory. This inefficiency limits the trainable model size and the available space for offloaded checkpoints, thereby constraining the maximum context length. Our work proposes several approaches to address the system memory bottleneck, enabling more cost-effective training for models with the desired size and context length.
@@ II Background and Related Works
II-A SSD offloading Systems
To overcome GPU memory limitations for large-scale model training, heterogeneous memory systems utilizing SSD offloading have emerged as a promising approach. These systems strategically move less frequently accessed data, primarily static model states, from expensive GPU memory to more abundant and cost-effective system memory or NVMe SSDs.
@@ II-A SSD offloading Systems
ZeRO-Infinity [21] represents a state-of-the-art SSD offloading system integrated within the DeepSpeed library [25]. It is based on the ZeRO distributed training strategy [4, 26], which partitions model states across data-parallel processes to eliminate memory redundancy. ZeRO-Infinity extends this by introducing the SSD offloading scheduler, enabling the offloading of partitioned model static states to CPU memory and subsequently to NVMe SSDs, as shown in Figure 1.
In the ZeRO-Infinity workflow, forward and backward passes are executed on the GPU, while the optimizer step is typically performed on the CPU since the compute complexity is lower than forward and backward, which is why it is not worth transferring it to the GPU for computation [26]. ZeRO-Infinity employs techniques such as bandwidth-centric partitioning and overlap-centric design to maximize data transfer efficiency across the PCIe bus and NVMe interfaces. While powerful for scaling model size, ZeRO-Infinity’s implementation, particularly its management of CPU system memory buffers for staging data between GPU and SSD, can introduce significant inefficiencies and overhead, especially on resource-limited machines with limited system memory addressed by this work.
@@
Figure 1: ZeRO-Infinity data flow during the backward pass for an n-layer model with three data-parallel ranks. Parameters P0(i) are moved to GPU, gradients G0(i) are processed and offloaded, and optimizer states A0(i) are updated on the CPU.
II-B Related Works in SSD offloading
II-B1 LoHan
LoHan [22] focuses specifically on enabling the fine-tuning of very large models on a single consumer-grade GPU within a commodity server that has limited main memory. It achieves this through holistic management of tensor movement, introducing techniques like active gradient offloading (overlapping CPU optimizer work with GPU backward pass) and traffic-aware activation swapping between GPU, CPU, and SSD. While LoHan demonstrates impressive capability in extreme resource constraints (e.g., fine-tuning 175B on one RTX 4090), its optimizations and design are primarily tailored for this single-GPU consumer hardware scenario. with brief discussion about multi-GPU scenario. Our work, MemAscend, addresses system memory inefficiencies within the general SSD offloading system like ZeRO-Infinity, which is designed for distributed training across various scales (single node to large clusters) and does not focus solely on the single-GPU or few GPUs use case. Therefore, LoHan addresses a different problem context.
II-B2 SSDTrain
@@
SSDTrain [27] proposes an adaptive system specifically for offloading activations to NVMe SSDs to mitigate the GPU memory bottleneck caused by large activation tensors, particularly during training with large batch sizes or long sequences. Specifically, it focuses on efficiently overlapping the I/O transfers of activations with GPU computation, demonstrating significant reductions in peak activation memory usage with negligible performance overhead. SSDTrain’s approach is complementary to systems like ZeRO-Infinity (and our MemAscend), which primarily handle the offloading of static model states. Techniques for activation offloading, like SSDTrain’s, can be potentially integrated with model state offloading systems to further optimize overall memory usage. As a result, SSDTrain addresses a different component of memory and is compatible with our offloading strategy.

### v2

@@ Abstract
Owing to the huge success of generative artificial intelligence (AI), large language models (LLMs) have emerged as a core subclass, powering applications such as question answering, text generation, and code completion. While fine-tuning these models on domain-specific data can yield significant performance gains, it also poses daunting computational challenges, especially for researchers and organizations with limited hardware resources. Although SSD offloading (i.e., ZeRO-Infinity) has emerged as a viable strategy to overcome the GPU memory barrier via leveraging both system memory (i.e., CPU DRAM) and storage space (i.e., solid-state devices, SSDs), its design primarily targets model-centric performance issues. As a result, key system-level issues, including system memory fragmentation, inefficient pinned buffer allocation, peak system memory usage spikes, and file system overhead, remain unaddressed, limiting scalability and inflating costs. Such an observation motivates this paper to introduce MemAscend, a framework that systematically tackles the underexplored system memory bottlenecks in SSD-offloaded LLM training, with a focus on resource-constrained environments. By streamlining pinned-memory allocation, eradicating fragmentation, and mitigating peak system memory usage, MemAscend reclaims a substantial system memory budget, enabling larger models, longer context windows, and higher batch sizes without exceeding modest hardware limits. Across diverse LLM benchmarks, MemAscend reduces peak system-memory consumption by an average of 55.7% compared with state-of-the-art SSD offloading techniques, lowering the hardware barrier for fine-tuning and unlocking new possibilities for cost-effective large-scale training on limited-resource machines.
@@ I Introduction
The growth of large-language-model (LLM) applications has significantly surged in recent years, propelled by steady advances in model capabilities [1]. LLMs have now powered conversational assistants, automatic summarization, multilingual translation, and countless other tasks [2, 3, 4, 5, 6]. To adapt LLMs to a specific domain, fine-tuning on domain-specific datasets is a widely used approach [7] to customize pre-trained foundational models and address the nuances of a specific domain better. Although fine-tuning markedly boosts reliability and task performance, it also demands considerable computation and memory usage. Commercial leaders routinely train foundation models on massive GPU clusters. For example, Meta’s Llama 3 relied on 16,000 NVIDIA H100 GPUs [3]. By contrast, academic labs, individual researchers, and small enterprises are often operating within tight hardware budgets, with limited GPU count and modest system-memory capacity. Finding cost-effective ways to fine-tune LLMs under such tight resource constraints, therefore, remains an open challenge.
@@
Broadly, LLM training consumes two memory classes: (1) static memory, which includes model weights and optimizer states, and (2) residual states, such as activations and intermediate values [8]. For residual states, activations dominate due to their linear growth with batch size and sequence length [9], and several techniques have been proposed to mitigate the above memory pressure. For instance, to lower activation memory, gradient checkpointing [10] is designed not to store every activation during the forward pass, but instead stores only a few checkpointed activations and recomputes those missed ones during backpropagation. The intermediate values, which are temporary results that persist across multiple operations, such as across multiple CUDA kernels [^1] CUDA (Compute Unified Device Architecture) kernels are functions, also basic execution unit, written in CUDA C/C++ that run on the GPU. [11], significantly contribute to peak memory usage. To lower the size of intermediate values, methods like Liger-Kernel [12] fuse operations (e.g., RMSNorm, SwiGLU, Cross-Entropy [13, 14, 15]) into single Triton [16] kernels, eliminating transient buffers and launch overhead. Additionally, memory-efficient kernels like Flash-Attention [17] tile the computation to avoid materializing the entire attention matrix at once, thereby reducing the size of intermediate values. Other approaches, such as mixed-precision training [18] and related quantization techniques, can reduce the memory footprint of both activations and intermediate values by casting tensors to fp16 or bf16.
@@
On the other hand, static memory is substantial because modern transformers contain billions of parameters, and optimizers such as Adam store both momentum and variance for each weight [19]. To lower the amount of static memory, quantization has been widely explored. For instance, 8-bit optimizers [20] and 4-bit optimizers [21] have shown promising results in reducing static memory consumption with minimal impact on convergence speed. On the other hand, Parameter-Efficient Fine-Tuning (PEFT) [22], especially Low-Rank Adaptation (LoRA) [23], significantly reduce the number of trainable parameters by focusing on low-rank matrices while keeping the pre-trained model frozen. Although methods like LoRA offer substantial computational and memory efficiency, they often require meticulous hyperparameter tuning, such as determining the rank and selecting which weights to train, to achieve optimal performance [24, 25]. Consequently, full-parameter training remains crucial for maximizing task-specific performance, particularly in scenarios where pre-trained models have not been exposed to domain-specific or private data. In such cases, the discrepancies between the target and pre-trained domains can be substantial, necessitating full-parameter training to bridge the gap effectively.
@@
Although full-parameter fine-tuning can achieve optimal performance in fine-tuning scenarios, the large size of pre-trained models limits the feasibility of training in resource-limited environments. Typically, fine-tuning tasks involve relatively smaller datasets, so the primary bottleneck in resource-limited scenarios is that GPU memory is insufficient to handle the static and residual memory requirements of these models. To address these limitations, SSD offloading has emerged as a promising solution. SSD offloading transfers static memory, such as model weights and optimizer states, from GPUs to SSDs, which generally provide significantly larger storage capacity. For example, ZeRO-Infinity [26], a state-of-the-art SSD offloading technique, moves static memory to NVMe SSDs. With SSD offloading, GPU memory is dedicated to storing intermediate values during computation.
System memory serves as an intermediary, either facilitating data transfers between the GPU and SSD or handling optimizer steps on the CPU to reduce data movement across the PCIe bus. Although SSD offloading may reduce training performance due to additional PCIe transfer latency, this impact is mitigated for fine-tuning tasks since fine-tuning tasks typically involve a relatively smaller dataset when compared with pre-trained tasks. Furthermore, SSD offloading, which targets alleviating the static memory bottleneck, can be integrated effectively with residual memory optimization techniques, such as Flash-Attention, Liger-Kernel, gradient checkpointing, or further offloading the checkpointed gradient into system memory. By integrating these techniques, GPU memory usage is further lowered, enabling the largest trainable context length or batch size, and potentially mitigating the original performance loss caused by additional data transfers across the GPU, system memory, and SSDs.
However, existing SSD offloading techniques (i.e., ZeRO-Infinity) are primarily designed to address model-centric performance issues. While GPU memory usage has been extensively studied, the system memory usage has received comparatively less attention and, according to our investigation, the system memory usage has emerged as a new bottleneck in the SSD offloading scenario. In this study, key inefficiencies are found through analyzing system memory requirements for SSD offloading. For example, although pinned memory [27] (i.e., cudaHostAlloc) can be used to facilitate GPU-to-CPU data transfers, its current management with ZeRO-Infinity causes system memory fragmentation that wastes 72.71% of buffer pool memory across different models, introducing unnecessary allocation overhead for system memory. Specifically, aggressive alignment in pinned memory allocation, while beneficial in data transfers, results in additional overhead and can double the allocation cost. Furthermore, inefficient CPU overflow checking contributes to peak memory overhead, leading to up to 1.25 times more memory usage depending on the size of the process tensor. In the end, our investigation reveals that these unnecessary system memory usage scales with model size, consistently causing an average of 55.7% system memory waste across different model sizes. This inefficiency limits the trainable model size and the available space for offloaded checkpoints, thereby constraining the maximum context length and the ability to use larger batch sizes, leading to suboptimal performance due to insufficient batch size. Although several works have been proposed after ZeRO-Infinity, such as [28, 29, 30], for further performance or efficiency enhancement, these approaches either focus on different aspects than ZeRO-Infinity or target specific scenarios, making them unsuitable for broader training use cases. As a result, the major unnecessary waste of system memory still remains unaddressed.
To address these challenges, this paper introduces MemAscend, a framework that aims to minimize system memory requirements within SSD offloaded fine-tuning while maximizing trainable model size and context length. MemAscend comprises four key components: (1) an adaptive buffer pool, (2) an alignment-free pinned memory allocation, (3) a fused overflow check mechanism, and (4) a direct NVMe engine. Firstly, the adaptive buffer pool eliminates memory fragmentation by adaptively fitting offloaded tensors into buffers on the system memory with the most suitable size. Secondly, the alignment-free pinned memory allocation avoided the alignment overhead between pinned memory allocations by implementing custom operations (i.e., c++ extensions) to avoid unnecessary alignment when allocating memory space. Thirdly, the fused overflow check mechanism is proposed to merge existing overflow-checking operations to eliminate the issue of double-peak memory usage. Lately, the direct NVMe engine bypasses the file system and enables direct data management on SSDs for low-latency SSD access. This study also demonstrates that the I/O transfer volume can be further reduced by integrating a pure half-precision optimizer, which transfers parameters, gradients, and momentum in half-precision at each optimizer step. Based on the above, the contributions of this study are summarized as follows:
@@
    Comprehensive system memory optimization: MemAscend reduces peak system memory consumption by an average of 55.7% through adaptive buffering, alignment-free pinned memory allocation, and removal of overflow-induced memory duplication.
@@
    Enhanced SSD I/O efficiency: By integrating a Direct NVMe Engine and a half-precision optimizer, MemAscend reduces I/O transfer volume by 58% and improves throughput by up to 24.21% or 56.80%, depending on the hardware configuration, while eliminating filesystem and staging overheads.
@@
    Improved scalability for fine-tuning scenario: On a system with 128 GiB of memory, MemAscend enables scaling context lengths from 16,384 to 131,072 tokens or increasing batch size from 4 to 32 under the same configuration.
@@ II Background and Related Works
This section provides background and related works by reviewing key developments in memory management for large-scale model training. It first outlines SSD offloading frameworks that extend GPU memory capacity using heterogeneous storage. It then discusses related SSD offloading techniques and their applicability across different training scenarios. Finally, it highlights GPU memory efficiency optimizations that complement offloading strategies.
II-A SSD offloading framework
@@
To address GPU memory limitations in large-scale model training, heterogeneous memory systems that incorporate SSD offloading have emerged as a promising solution. These systems strategically relocate less frequently accessed data, primarily static model states, from expensive GPU memory to more abundant and cost-effective system memory or NVMe SSDs. ZeRO-Infinity [26] is a state-of-the-art (SOTA) SSD offloading system integrated into the DeepSpeed library [31]. Building on the ZeRO distributed training strategy [8, 32], which partitions model states across data-parallel processes to eliminate memory redundancy, ZeRO-Infinity extends the design by introducing an SSD offloading scheduler. This scheduler enables partitioned static model states to be offloaded first to CPU memory and subsequently to NVMe SSDs, as illustrated in Figure 1.
@@
Figure 1: ZeRO-Infinity data flow during the backward pass for an $n$-layer model with three data-parallel ranks. In this example, network refers to the use of peer-to-peer communication operations, such as Allgather or ReduceScatter, which synchronize data (i.e., tensors) across multiple GPUs during distributed training. Parameters P0(i), which are the first tensors for the i-th GPU, are moved to each GPU. The gradients G0(i), which correspond to P0(i), are processed and offloaded. The optimizer states O0(i), which correspond to P0(i) and G0(i), are updated on the CPU.
In ZeRO-Infinity’s workflow, forward and backward passes are executed on the GPU, while the optimizer step is typically handled on the CPU due to its lower computational complexity. This is because transferring the optimizer step to the GPU is often not worthwhile, given the overhead involved [32]. To optimize data movement, ZeRO-Infinity employs techniques such as bandwidth-centric partitioning and overlap-centric design, which aim to improve transfer efficiency across the PCIe and NVMe interfaces. While bandwidth-centric partitioning partitions one large tensor and stagger multiple tensors to fully saturate available NVMe bandwidth during each transfer window, the overlap-centric design asynchronously loads the next set of weights or gradients from SSDs into system memory during forward passes executing on GPUs. Despite its effectiveness in scaling model size, ZeRO-Infinity’s reliance on CPU system memory as a staging buffer between the GPU and SSD can lead to significant inefficiencies. These overheads become particularly pronounced on resource-constrained systems with limited system memory, an issue that this work seeks to address.
II-B Related Works of SSD offloading
Following the introduction of the state-of-the-art SSD offloading system, ZeRO-Infinity, subsequent studies have investigated various strategies within the broader context of SSD offloading. While some of these approaches, such as activation offloading, are compatible with ZeRO-Infinity and can be integrated to further enhance its performance, others are designed for specific hardware setups or constrained training environments, addressing distinct use cases that complement ZeRO-Infinity but do not directly generalize to its distributed training framework. This section reviews these related works in detail.
II-B1 SSDTrain
@@
SSDTrain [30] is an adaptive system designed to offload activations to NVMe SSDs, mitigating GPU memory bottlenecks caused by large activation tensors, particularly in training workloads with large batch sizes or long sequences. The system emphasizes efficient overlap between I/O transfers and GPU computation, achieving substantial reductions in peak activation memory usage with negligible performance overhead. SSDTrain’s focus on activations complements systems like ZeRO-Infinity (and our MemAscend), which primarily offload static model states. Activation offloading techniques, such as those in SSDTrain, can potentially be integrated with model state offloading systems to further optimize end-to-end memory usage. As such, SSDTrain targets a different memory component and remains compatible with our offloading strategy.
II-B2 Smart-Infinity
Smart-Infinity [29] tackles the storage bandwidth bottleneck in SSD-offloaded LLM training by utilizing near-storage processing (NSP) hardware. It introduces SmartUpdate, a mechanism that offloads parameter updates to custom accelerators within computational storage devices (CSDs), thereby reducing data movement between storage and system memory. In addition, Smart-Infinity features an efficient data transfer handler to overlap communication with fixed memory usage and employs accelerator-assisted gradient compression and decompression to improve scalability. While these innovations significantly accelerate LLM training, Smart-Infinity relies on specialized hardware, limiting its applicability to systems equipped with NSP-enabled storage. In contrast, MemAscend addresses system memory inefficiencies in general SSD offloading systems like ZeRO-Infinity, which support distributed training from single-node setups to large-scale clusters without dedicated hardware. Thus, Smart-Infinity and MemAscend target different optimization contexts.
II-B3 LoHan
LoHan [28] is tailored for fine-tuning very large models on a single consumer-grade GPU within a commodity server with limited main memory. It achieves this through holistic tensor management, introducing techniques such as active gradient offloading, which overlaps CPU-based optimizer execution with the GPU’s backward pass, and traffic-aware activation swapping across GPU, CPU, and SSD tiers. LoHan demonstrates impressive capabilities under extreme resource constraints (e.g., fine-tuning a 175B model on a single RTX 4090). However, its design is primarily optimized for single-GPU environments, with only brief consideration for multi-GPU setups. In contrast, MemAscend targets system memory inefficiencies in SSD offloading systems like ZeRO-Infinity, designed for distributed training across a broader range of hardware scales. Therefore, LoHan and MemAscend serve distinct use cases within the LLM training optimization.